OpenStudy (anonymous):

HELP! WILL MEDAL AND FAN!! Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (ybarrap):

Add the 1st equation to the negative of the 2nd equation. What do you get?

OpenStudy (anonymous):

Y= -2x + 1 and Y= -4x - 1

OpenStudy (anonymous):

Y= -2x + 1 and Y= -4x - 1

OpenStudy (ybarrap):

Like this -2x + y = 1 plus -(-4x + y = -1) equals -2x -(-4x) + y - y = 1 -(-1) Do you see what happened here?

OpenStudy (anonymous):

yeah

OpenStudy (ybarrap):

So what do you think we should do with that last equation?

OpenStudy (anonymous):

solve it or simplify it

OpenStudy (ybarrap):

Ok, try it out - I'll watch and help

OpenStudy (anonymous):

are you tring to put it into y= form?

OpenStudy (ybarrap):

How can we simplify this - -2x -(-4x) + y - y = 1 -(-1) ?

OpenStudy (anonymous):

-2x+y=1

OpenStudy (ybarrap):

See the y - y will cancel to become zero $$ -2x -(-4x) + 0 = 1 -(-1)\\ $$ What's else can we do?

OpenStudy (anonymous):

subtract the values and get -6x

OpenStudy (ybarrap):

Close The -(-4x) becomes just 4x: $$ -2x +4x + 0 = 1 -(-1)\\ $$ Now what?

OpenStudy (ybarrap):

What is -2x + 4x ?

OpenStudy (anonymous):

2x

OpenStudy (ybarrap):

yes! So now we have $$ 2x+0=1-(-1) $$ What's can we do next?

OpenStudy (anonymous):

2x+0=-2

OpenStudy (ybarrap):

-(-1) = 1 Remember, negative of a negative number makes it positive Try again

OpenStudy (anonymous):

2x+0=1

OpenStudy (ybarrap):

$$ 2x+0=1-(-1)=1+1 $$ Do you see that? What's next?

OpenStudy (anonymous):

2x+0=2

OpenStudy (ybarrap):

yes! Ok, what do we do now?

OpenStudy (anonymous):

subtract 0 from both sides and get 2x=2

OpenStudy (ybarrap):

What next?

OpenStudy (anonymous):

the answer choices are A. (3,1) B (-1,3) C (-1,-3) D (1,3)

OpenStudy (ybarrap):

We need x by itself, what do we need to divide out on both sides to make 2x equal to just x ?

OpenStudy (anonymous):

2 divided by both sides

OpenStudy (ybarrap):

Yes! We need to divide 2 to both sides: $$ \cfrac{2x}{2}=\cfrac{2}{2}=? $$ What does this give us for x?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

x=1

OpenStudy (ybarrap):

Yes! Now you know one variable. Now let's use one of your original equations to get y Let's use this one $$ -2x + y = 1 $$ Plug in x=1 and see what you get for y.

OpenStudy (ybarrap):

$$ -2\times 1 + y = 1 $$ So now what is y?

OpenStudy (anonymous):

-2y=1

OpenStudy (ybarrap):

$$ -2\times 1 + y = -2+y=1 $$ See that?

OpenStudy (anonymous):

yeah

OpenStudy (ybarrap):

How do you move the -2 to the other side of the equal sign?

OpenStudy (anonymous):

im so confused this isnt leading to any of the answer choices

OpenStudy (anonymous):

and i have so many more questions in so little time

OpenStudy (ybarrap):

Add 2 to both sides of this equation $$ -2+y+2=1+2 $$ What does this give us?

OpenStudy (anonymous):

-2 + 2y = 3

OpenStudy (ybarrap):

Almost there $$ -2+2+y=3 $$

OpenStudy (ybarrap):

What's -2 + 2?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

y=3

OpenStudy (ybarrap):

Yes! So now you know everything. x=1 y=3 (1,3)

OpenStudy (anonymous):

oh thats make so much sense thanks so much

OpenStudy (ybarrap):

When you do one on your own, all the others will be like adding 1+1=2

OpenStudy (ybarrap):

do another on your own right here

OpenStudy (ybarrap):

Have some courage my friend, everything takes time -- don't worry about the little things.